OpenStudy (anonymous):

A goods train 158 metres long, leaves Delhi at 6 a.m and travels at an average speed of 32 km/hr. A mail train 130 metres long, leaves Delhi at 12 noon and follows the goods train travelling at an average speed of 80 km/hr. At what time and at what distance from Delhi will the mail train completely cross the goods train?

OpenStudy (anonymous):

length1=158,speed 32km/hr.length2=130,time1=x,time2=x-6,speed2=80km/hr

OpenStudy (anonymous):

32*x=80*(x-6)

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

@anteater

OpenStudy (anonymous):

@anteater ,hahha

OpenStudy (solomonzelman):

158 meters long :)

OpenStudy (anteater):

LOL ... I am thinking about it!

OpenStudy (anteater):

And the other guy is replying, so I will continue on paper and verify what he says. :)

OpenStudy (solomonzelman):

Okay, so the first train is 158meter, 32 mph. LEAVES at 6:00 am the second train is 130meter, 80mph. LEAVES at 12 noon (6 hours after the 1st train). (All the above is referring to where the very back of each train is) You know that the 1st train is 6 hours ahead, which is in distance `32 × 6` and equals 192 miles. So the first train is 192 miles ahead. So now they are traveling the same direction and when will the second train (the one that travels faster) be equal (equal with fronts, not equal with the backs of the trains)? Let x be number of hours. 192 + 32x = 80x 192 =48x x=4 So thus the trains' backs will be at the same distance (from Delhi) after 4 hours. Keep in mind though, that the firs train is 158 meters long, and the second is 130 meters long. All you have left is to calculate the short time during which the second train is getting from being even by the backs, to being even by the fronts.

OpenStudy (solomonzelman):

Let me know if you need help calculating that.

OpenStudy (anteater):

Actually, since it says the second train has to clear the first train, I think they mean the back of the second train will have to be even with the front of the first train. Correct me if I am mistaken, please.

OpenStudy (solomonzelman):

Well.... yeah, perhaps the backs have to be even. Then it is simpler :)

OpenStudy (anteater):

No, I didn't mean the backs have to be even. :) I meant that the back of the second train has to be even with the nose of the first train, since it wants to know at what distance it will "completely cross" the goods train.

OpenStudy (anonymous):

srry,322 meters 192+130

OpenStudy (anteater):

So by "completely cross", do they mean it is even with the first train (nose and nose) or entirely ahead of the first train?

OpenStudy (solomonzelman):

well.... sounds like the first train has to be completely behind the second one. I can't say this for sure though.

OpenStudy (anteater):

I wonder if they give a diagram with this problem so we could know for sure what they meant. dg2, do they have a picture of the trains?

OpenStudy (anonymous):

nope @anteater

OpenStudy (anteater):

By any chance do they give the answer so that you can verify the approach?

OpenStudy (anteater):

You could work it out both ways, I suppose, and tell your instructor that since the directions weren't clear you did it both ways. :)

OpenStudy (anonymous):

k tell me

OpenStudy (anteater):

So you could figure it one way, where the fronts of the trains are even. And then, you could do it another way where the back of the second train is even with the front of the first train. Extra work, I know ... but maybe they would give you extra credit? Never hurts to ask! :D

OpenStudy (anteater):

Sorry ... had to step away for a moment. Please let me know what you got for the nose-to-nose trains and/or what you got for the scenario where the second train is ahead of the first. I can double check. :)

OpenStudy (anonymous):

oh i dont know how yo do that @anteater

OpenStudy (anteater):

Just a moment, please :)

OpenStudy (anonymous):

this sum is boring :-)

OpenStudy (anteater):

Sorry!

OpenStudy (anteater):

Working on a drawing

OpenStudy (anonymous):

k

OpenStudy (anteater):

I think they want the back of the second train to be even with the front of the first train, really.

OpenStudy (anteater):

The more I think about it, the more I am sure that is what they want.

OpenStudy (anteater):

So ... just a moment, please.

OpenStudy (anonymous):

we can close this problem .later we do.shall we move on to next prob

OpenStudy (anteater):



OpenStudy (anteater):

Sure - if you want, I can work on this one and email you later?

OpenStudy (anteater):

And we can compare

OpenStudy (anonymous):

yea!3q frnd

OpenStudy (anonymous):

then.......

OpenStudy (anonymous):

shall we move to next

OpenStudy (anteater):

Sure :)

OpenStudy (anonymous):

do u know simple interest

OpenStudy (anteater):

Yes P = rt

OpenStudy (anteater):

Oops

OpenStudy (anteater):

I = Prt ... don't know what I was thinking there!

OpenStudy (anonymous):

Robin lends Rs.9 to Rahul on the condition that the loan is repayable in 10 months in 10 equal installments of Re.1 each. Find the rate of interest per annum.

OpenStudy (anonymous):

1=9*10*r/100 r=100/90=10/9

OpenStudy (anonymous):

simple interest=pricipal*rate *time/100

OpenStudy (anonymous):

is it correct?

OpenStudy (anteater):

Yes, that should be correct.

OpenStudy (anonymous):

but in options 22*1/3,25*2/3,26*2/3,28*1/3

OpenStudy (anteater):

Give me a moment to think on this one. They have thrown a twist in here by making the loan due in 10 months, rather than 1 year.

OpenStudy (anteater):

I think we could work backward on it by starting with the payment. They say that you pay back 1 Re. each in 10 installments in 10 months.

OpenStudy (anteater):

So, how much money do you pay altogether?

OpenStudy (anonymous):

9*10*r/100*12

OpenStudy (anonymous):

total rs10

OpenStudy (anteater):

Right. But the principal amount was Re.9, so how much total interest was paid?

OpenStudy (anonymous):

1

OpenStudy (anteater):

So you paid 1/9 of what you borrowed in 10 months. I think we could set up a ratio to solve it.

OpenStudy (anteater):

1/9/ 10 = x/12

OpenStudy (anteater):

Nope! I am thinking wrong on this.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we can ask help with @ganeshie8

OpenStudy (anteater):

I think that would be a very good idea! :D

ganeshie8 (ganeshie8):

are u guys working on this problem : ` Robin lends Rs.9 to Rahul on the condition that the loan is repayable in 10 months in 10 equal installments of Re.1 each. Find the rate of interest per annum.` ?

OpenStudy (anteater):

Yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

p+p*r*t/100*12

OpenStudy (anonymous):

annual instalment formula=100*p/100*n+n*(n-1)/2

OpenStudy (anteater):

Thank you!

OpenStudy (anonymous):

but didnt get the answer

OpenStudy (anteater):

Hmm.

OpenStudy (anteater):

I have a question, though ... would that be annually compounded interest or simple interest for that formula?

ganeshie8 (ganeshie8):

compounded every month on the balance amount

ganeshie8 (ganeshie8):

at the end of first month, balance = 9 + (23.5% of 9) Robin pays 1 rupee, so balance reduces to : 8 + (23.5% of 9)

ganeshie8 (ganeshie8):

this cycle repeats every month

OpenStudy (anteater):

That makes sense.